# GALLa: Graph Aligned Large Language Models for Improved Source Code Understanding

### v1

@@ Abstract
Programming languages possess rich semantic information such as data flow that is represented by graphs and not available from the surface form of source code. Recent code language models have scaled to billions of parameters, but model source code solely as text tokens while ignoring any other structural information. Conversely, models that do encode structural information of code make modifications to the Transformer architecture, limiting their scale and compatibility with pretrained LLMs. In this work, we take the best of both worlds with GALLa - Graph Aligned Large Language Model. GALLa utilizes graph neural networks and cross-modal alignment technologies to inject the structural information of code into LLMs as an auxiliary task during finetuning. This framework is both model-agnostic and task-agnostic, as it can be applied to any code LLM for any code downstream task, and requires the structural graph data only at training time from a corpus unrelated to the finetuning data, while incurring no cost at inference time over the baseline LLM. Experiments on five code tasks with four different baseline LLMs ranging in size from 350M to 8B validate the effectiveness of GALLa, demonstrating consistent improvement over the baseline, even for powerful models such as LLaMA3.
@@ 1 Introduction
In recent years, applying large language models (LLMs) to processing and generating source code has been a research topic of special interest in both natural language processing and software engineering community (Chen et al., 2021). However, unlike natural languages, programming languages have rich semantics besides the lexical representation of source code, such as the path of execution, the flow of data, and the relation of function calling. These semantics are represented by graph structures such as Abstract Syntax Tree (AST), Control Flow Graph (CFG), and Data Flow Graph (DFG). While enhancing LLMs with lexical representations of source code has the potential to boost their performance, the integration of these richer semantic constructs has yet to be fully realized (Zhang et al., 2024).
@@
Through extensive experiments on five code understanding and generation tasks, we validate the effectiveness of GALLa on four distinct base LLMs ranging in size from 350M to 8B. GALLa brings consistent improvement over all baseline models, and even demonstrates abilities to generalize structural knowledge acquired during graph alignment to programming languages that are absent in the alignment data. All the data used in our experiments (both graph alignment and downstream tasks) are sourced from publicly available datasets, and the complete code for reproducing our results will be released in https://github.com/codefuse-ai.
@@ 2 Related Work
The first category is represented by GraphCodeBERT (Guo et al., 2021), which finetunes CodeBERT (Feng et al., 2020) on concatenated source code and data flow graphs. Attention masks in the model are modified to reflect the graph structures: a node in the graph is allowed to attend another node , only if there is an edge from to , while a node and a source code token can attend to each other only if they correspond to each other in the graph. StructCoder (Tipirneni et al., 2024) also modifies attention masks similarly to encode the relations between source code tokens and AST, DFG tokens.
For the second category, TreeBERT (Jiang et al., 2021) is an encoder-decoder model where each input token to the encoder is a node’s constituent path in the AST, while the decoder is trained to generate the source code with cross-attention to the encoder representations. Many other works in this category, including SynCoBERT (Wang et al., 2021), SPT-Code (Niu et al., 2022), and UniXcoder (Guo et al., 2022) first map ASTs into text sequences (for example, by depth-first traversing), and then send them into code language models. However, this method only applies to simple graph structures such as AST, but not more complex ones such as DFG which may include loops.
@@
Despite their contributions, all these approaches focus primarily on the encoder of the Transformer and are not fully compatible with decoder-only LLMs. For instance, in the first group, the presence of cycles in the graphs means that the corresponding attention mask cannot retain a lower-triangular format, as required by LLMs. Adapting LLMs to these graph-structure-based attention masks would necessitate extensive retraining due to the mask’s inconsistency with the original causal language modeling objectives. In contrast, the proposed GALLa method processes graph information externally, allowing LLMs to remain unmodified.
Apart from these graph-enhanced models, there is another model TransCoder-IR (Szafraniec et al., 2023), which is specifically proposed for the code translation task and utilizes LLVM intermediate representation (IR) to ground the model’s understanding of code by generating source code from IR and vice versa. Like our method, TransCoder-IR only uses IR for alignment at training time but does not need it at test time. However, intermediate representations are represented in the same modality as the source code (i.e. text modality) and provide a different type of information compared with structural graphs.
@@ 3 Method
The objective of GALLa is to align language models to the implicit structures of code (represented by AST and DFG - see more details in Appendix A) when finetuning on code-related tasks, and an overview of the method is provided in Figure 1. The model consists of three modules: GNN encoder, adapter, and LLM decoder (Section 3.1). The training of GALLa includes two stages: graph encoder pretraining and graph-LLM alignment (Section 3.2 and Figure 2). The notations used in this section are presented in Table 1.
| a graph | |
| a node in the graph | |
| an edge in the graph | |
| number of nodes in the graph | |
| number of edges in the graph | |
| node feature dimension of the graph data | |
| GNN hidden and output dimension | |
| LLM hidden and embedding dimension | |
| number of graph tokens in LLM’s input | |
| number of text tokens in LLM’s input | |
| node features (input to GNN) | |
| edge indices (input to GNN) | |
| contextual node features (output of GNN) | |
| query vectors (input to adapter) | |
| embedding vectors (input to LLM) | |
| logits (output of LLM) |
@@ 3.1.1 GNN Encoder
To fully capture the rich information contained in the structural graphs such as loops, node degrees, and edge directions, we first process the graphs with a graph neural network (GNN) to extract node information. For a graph with nodes and edges, the GNN takes as input a node feature matrix and an edge matrix , where each element is a node index. The output of the GNN is the contextual node representations .
Since the graph data extracted from the code consists of only a list of nodes and the edge indices, we need to first encode each node’s corresponding code segment into a feature vector. This can be done using any code embedding model such as CodeT5+ (Wang et al., 2023). The GNN can be any convolution-based or self-attention-based directed GNN such as MagNet (Zhang et al., 2021) or DUPLEX (Ke et al., 2024), as DFG is a type of directed graph.
@@ 3.1.2 Adapter
The outputs of GNN are projected by an adapter into the LLM’s embedding space. Following Qwen-VL (Bai et al., 2023), we use a single layer of cross-attention as the adapter, where the GNN’s outputs serve as keys and values, while the queries are learnable vectors:
@@ 4 Experiments
In this section, we first discuss the details of our experiments in subsections 4.1 and 4.2, and then provide the results in subsections 4.3 and 4.4.
@@ 4.1.1 Graph Alignment
For graph alignment, we used 240K Python programs and 75K Java programs from CodeNet (Puri et al., 2021). For each program, we extracted one AST and one DFG using Program Structure Interface 222https://plugins.jetbrains.com/docs/intellij/psi.html, which resulted in 626K code-graph pairs after removing empty graphs and source code files that are longer than 4096 tokens.
To improve the generalization capability of our model and to align the representations of different programming languages for certain cross-lingual tasks such as code translation, we used a special type of AST and DFG - universal AST (UAST) and universal DFG (UDFG), where different languages share the same set of node types to the possible extent (see more details in Appendix A). In total, there are 43 node types in our graph data.
For the Graph2Code task, we removed all Python programs where no main function can be found (which consist more than 80% of the 240K Python programs), resulting in 150K Java graph-code pairs and 81K Python graph-code pairs.
@@
Edge prediction: given the graph tokens and the source code of two nodes, the model is tasked to predict where there is an edge between them. This task is constructed only on DFG data.
Parent prediction: given the graph tokens and the source code of one node, the model is tasked to predict the node’s parent in the graph. This task is constructed on both AST and DFG data.
Child prediction: given the graph tokens and the source code of one node, the model is tasked to predict the node’s children in the graph. This task is constructed on both AST and DFG data.
In our preliminary experiments, we found edge prediction data construct from AST graphs to bring no improvement, as connected nodes in the AST often have high textual overlap and can be trivially predicted. We also filtered out several other types of questions, including node classification (which can be easily done by only looking at the node’s source code) and some tasks that involve counting such as counting the number of nodes in the graph, the number of edges in the graph, and the number of node types in the graph, which prove to be too difficult for the model.
We wrote about ten question templates and ten answer templates for each type of question. When constructing the data, one question template and one answer template were randomly chosen for each sample, as diverse question templates have been found to improve cross-task generalization capability in instruction finetuning (Sanh et al., 2022). We sampled about 75K graph-question-answer tuples for each (language, graph type, question type) combination, which resulted in a total of 770K GraphQA data.
@@ 4.1.2 Downstream Tasks
In the main experiments, we consider the setting of multi-task finetuning (MFT), where the model is simultaneously trained on all five downstream tasks. However, we demonstrate with a small-scale experiment that GALLa can be also used for single-task fientuning (SFT).
Model Setting Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect Avg CodeGen 350M Baseline 40.2/42.3 94.4 9.0 13.6/8.8/12.5 56.9 34.7 G2C 42.7/43.3 94.6 8.5 13.9/9.5/12.8∗ 56.5 35.2 (+1%) G2C + GraphQA 50.0/45.3 93.9 8.6 14.0/9.5/12.7∗ 58.7 36.6 (+5%) StarCoder 1B Baseline 0.6/0.0 40.3 0.4 7.8/3.2/5.1 54.2 14.0 G2C 6.1/1.2 50.4 0.6 9.4/4.3/6.8 54.8 16.7 (+20%) G2C + GraphQA 6.1/1.2 65.8 0.3 9.5/4.5/7.4 56.5 18.9 (+36%) Phi-1 1.3B Baseline 73.8/53.3 93.5 10.1 14.6/11.3/13.8 61.1 41.4 G2C 67.7/64.3 94.3 10.9 15.1/11.1/13.4∗ 60.4 42.2 (+2%) G2C + GraphQA 72.0/66.3 94.9 10.2 15.2/11.7/13.6∗ 60.3 43.0 (+4%) LLaMA3 8B Baseline 80.5/74.8 94.7 12.1 14.2/11.2/12.3 56.7 44.6 G2C 80.5/77.0 94.9 13.4 14.0∗/11.7/11.8 56.7 45.0 (+1%) G2C + GraphQA 80.5∗/78.2 94.9 13.6 14.5/11.7/12.9 57.1 45.4 (+2%)
@@ 4.2.1 Model
We use a DUPLEX (Ke et al., 2024) with 1024 hidden states and 7M parameters as the GNN, and a cross-attention layer with randomly initialized learnable queries as the adapter. For the LLM, we consider four distinct models of varying sizes: CodeGen 350M (Nijkamp et al., 2023), StarCoder 1B (Li et al., 2023), Phi-1 1.3B (Gunasekar et al., 2023), and LLaMA3 8B (Dubey et al., 2024). All of these models are prerained (partially) on source code data, and demonstrate strong performance on code-related downstream tasks.
@@ 4.2.2 Training
The second stage of training largely follows the same setting but with a smaller learning rate (5e-5) and a smaller global batch size (96). The model is trained for 5 epochs on the mixture of downstream task data and graph alignment data in stage 2, and the checkpoint with the lowest validation loss on the downstream tasks is chosen for evaluation. Among the four LLMs, CodeGen, StarCoder, and Phi-1 are trained in full scale, while LLaMA3 is trained using LoRA (Hu et al., 2022) with rank 64. All training in this stage takes place on a single machine with 8 A100s.
@@ 4.3 Results
The results of applying Graph2Code and GraphQA alignment to the code finetuning process of four models are presented in Table 3. Graph alignment brings consistent improvement over the baseline model, increasing the average performance on five tasks by up to 36%. Comparing the results between the four models, we find that GALLa’s effectiveness is most prominent on the weakest baseline StarCoder, while relatively less notable on the strongest model LLaMA3.
To verify the statistical significance of these improvements, we conducted Chi squared tests on the tasks with discrete performance metrics (i.e. code translation, code repair, clone detection, and defect detection) and Wilcoxon signed-rank test on the tasks with continuous performance metrics (i.e. code summarization). Most of the differences are significant, except for CodeGen and Phi-1 on JavaScript summarization, and LLaMA3 on Java-to-Python translation and Python summarization. The complete table of p values is given in Table 10 in Appendix B.
Notably, while our graph alignment data include only Python and Java programs, from Table 3 we observe that they can even improve three of the four models’ performance on tasks in other languages - code summarization in JavaScript, and defect detection in C. This showcases that the knowledge about code structures acquired in GALLa can be generalized across programming languages, as learning to align to Python and Java structural graphs improves the finetuning performance of downstream tasks in other languages.
In Table 4, we also present the results of single-task finetuning with Phi-1. In this setting, we find that Graph2Code still improves the average performance on all tasks by 2%, and by 3% on Python and Java tasks. On the other hand, GraphQA brings limited improvement, and even a large drop on the C defect detection task. We hypothesize that this is because the diverse data in GraphQA serve a similar purpose to instruction tuning for LLMs (Sanh et al., 2022; Chung et al., 2022), where the benefits of cross-task transfer only start to manifest when the number of tasks is large.
@@
| Baseline | 73.8/53.3 | 93.5 | 10.1 | 14.6/11.3/13.8 | 61.1 | 41.4 |
| G2C | 73.2/66.9 | 95.2 | 11.7 | 15.1/11.6/14.1 | 61.6 | 43.7 |
| G2C + GQA | 76.2/68.9 | 95.3 | 12.7 | 14.9/12.1/13.6 | 59.9 | 44.2 |

### v2

@@ Abstract
Programming languages possess rich semantic information - such as data flow - that is represented by graphs and not available from the surface form of source code. Recent code language models have scaled to billions of parameters, but model source code solely as text tokens while ignoring any other structural information. Conversely, models that do encode structural information of code make modifications to the Transformer architecture, limiting their scale and compatibility with pretrained LLMs. In this work, we take the best of both worlds with GALLa - Graph Aligned Large Language Models. GALLa utilizes graph neural networks and cross-modal alignment technologies to inject the structural information of code into LLMs as an auxiliary task during finetuning. This framework is both model-agnostic and task-agnostic, as it can be applied to any code LLM for any code downstream task, and requires the structural graph data only at training time from a corpus unrelated to the finetuning data, while incurring no cost at inference time over the baseline LLM. Experiments on five code tasks with seven different baseline LLMs ranging in size from 350M to 14B validate the effectiveness of GALLa, demonstrating consistent improvement over the baseline, even for powerful models such as LLaMA3 and Qwen2.5-Coder.
@@ 1 Introduction
In recent years, applying large language models (LLMs) to processing and generating source code has been a research topic of special interest in both natural language processing and software engineering community (Chen et al., 2021). However, unlike natural languages, programming languages have rich semantics besides the lexical representation of source code, such as the path of execution, the flow of data, and the dependency of function calling. These semantics are represented by graph structures such as Abstract Syntax Tree (AST), Control Flow Graph (CFG), and Data Flow Graph (DFG). While enhancing LLMs with lexical representations of source code has the potential to boost their performance, the integration of these richer semantic constructs has yet to be fully realized (Zhang et al., 2024).
@@
Through extensive experiments on five code understanding and generation tasks, we validate the effectiveness of GALLa on seven distinct base LLMs ranging in size from 350M to 14B. GALLa brings consistent improvement over all baseline models, and even demonstrates abilities to generalize structural knowledge acquired during graph alignment to programming languages that are absent in the alignment data. All the data used in our experiments (both graph alignment and downstream tasks) are sourced from publicly available datasets111https://huggingface.co/datasets/codefuse-ai/GALLa, and the complete code for reproducing our results is released in https://github.com/codefuse-ai/GALLa.
@@ 2 Related Work
The first category is represented by GraphCodeBERT (Guo et al., 2021), which finetunes CodeBERT (Feng et al., 2020) on concatenated source code and DFGs. Attention masks in the model are modified to reflect the graph structures: a node token is allowed to attend another node token , only if there is an edge from to in the graph, while a node and a source code token can attend to each other only if they correspond to each other. StructCoder (Tipirneni et al., 2024) also modifies attention masks similarly to encode the relations between source code tokens and AST, DFG tokens.
For the second category, TreeBERT (Jiang et al., 2021) is an encoder-decoder model where each input token to the encoder is a node’s constituent path in the AST, while the decoder is trained to generate the source code. Many other works in this category, including SynCoBERT (Wang et al., 2021), SPT-Code (Niu et al., 2022), and UniXcoder (Guo et al., 2022), map ASTs into text sequences (for example, by depth-first traversing) as model input. However, this method only applies to simple graph structures such as AST, but not more complex ones such as DFG, which may include loops.
@@
Despite their contributions, all these approaches focus primarily on the encoder of the Transformer and are not fully compatible with decoder-only LLMs. For instance, in the first group, the presence of cycles in the graphs means that the corresponding attention mask cannot retain a lower-triangular format, as required by LLMs. Adapting LLMs to these graph-structure-based attention masks would necessitate extensive retraining due to the mask’s inconsistency with the original causal language modeling objectives. In contrast, the proposed GALLa method processes graph information externally, allowing the LLM architecture to remain unmodified.
Apart from these graph-enhanced models, there is another model TransCoder-IR (Szafraniec et al., 2023), which utilizes LLVM intermediate representation (IR) to ground the model’s understanding of code by generating source code from IR and vice versa. Like our method, TransCoder-IR only uses IR for alignment at training time but does not need it at test time. However, intermediate representations are also text tokens and provide limited structural information.
@@ 3 Method
The objective of GALLa is to align language models to the implicit structures of code (represented by AST and DFG - see more details in Appendix A) when finetuning on code-related tasks, and an overview of the method is provided in Figure 1. In this section, we elaborate on two key challenges: 1) how to input graph information into LLMs, and 2) how to design training tasks so that LLMs can learn about such graph information. For the first challenge, we design a model that consists of three modules: GNN encoder, adapter, and LLM decoder (Section 3.1). For the second challenge, we propose a two-stage training scheme: graph encoder pretraining and graph-LLM alignment (Section 3.2 and Figure 2). The notations used in this section are presented in Table 1.
a graph a node in the graph an edge in the graph number of nodes in the graph number of edges in the graph node feature dimension of the graph data GNN hidden and output dimension LLM hidden and embedding dimension number of graph tokens in LLM’s input number of text tokens in LLM’s input node features (input to GNN, extracted by a text encoder) edge indices (input to GNN) contextual node features (output of GNN) query vectors (input to adapter) embedding vectors (input to LLM) logits (output of LLM)
@@ 3.1.1 GNN Encoder
To fully capture the rich information contained in the structural graphs such as loops, node degrees, and edge directions, we first process the graphs with a graph neural network (GNN) to extract node information. For a graph with nodes and edges, a text encoder is used to extract node features from the corresponding source code of each node. These features are used to construct the node feature matrix , which are fed into the GNN along with the edge matrix , where each element is a node index. The output of the GNN is the contextual node representations .
@@ 3.1.2 Adapter
The outputs of GNN are projected by an adapter into the LLM’s embedding space. Following Qwen-VL (Bai et al., 2023), we use a single layer of cross-attention as the adapter, where the GNN’s outputs serve as keys and values, and the queries are learnable vectors:
@@ 3.2 Training Procedure
Motivated by the pretraining + instruction finetuning paradigm in LLMs, we divide the training of GALLa into two stages. The first stage is self-supervised (continue) pretraining, where only AST/DFG data extracted from source code are used. The second stage is supervised finetuning, where GraphQA data are collected by designing graph-related questions and manually crafting question-answer templates.
@@ 3.2.2 Stage 2: Graph-LLM Alignment
3.2.3 Inference
After aligning the source code representation of LLMs to code graph structures in training stage 2, at inference time the graph encoder and adapter are discarded, so that the LLM can respond to user queries using its internal knowledge without any loss of speed compared with the base LLM. We choose not to encode the code in user queries with GNN at inference time because 1) it would require extracting AST and DFG from the code first, resulting in an unreasonable inference latency; 2) extracting AST and DFG requires that the input code is syntactically correct and complete, which is not true for many downstream tasks such as defect detection and code repair.
@@ 4 Experiments
In this section, we first discuss the details of our experiments (4.1, 4.2), and then provide the results in 4.3 and 4.4.
@@ 4.1.1 Graph Alignment
For graph alignment, we used 240K Python programs and 75K Java programs from CodeNet (Puri et al., 2021). For each program, we extracted one AST and one DFG using Program Structure Interface 333https://plugins.jetbrains.com/docs/intellij/psi.html, which resulted in 626K code-graph pairs after removing empty graphs and source code files that are longer than 4096 tokens. More details about these AST and DFG can be found in Appendix A.
For the Graph2Code task, we removed all Python programs where no main function can be found, resulting in 150K Java graph-code pairs and 81K Python graph-code pairs.
@@
1) Edge prediction: given the graph tokens and the source code of two nodes, the model is tasked to predict where there is an edge between them. This task is constructed only on DFG data.
2) Parent prediction: given the graph tokens and the source code of one node, the model is tasked to predict the node’s parent in the graph. This task is constructed on both AST and DFG data.
3) Child prediction: given the graph tokens and the source code of one node, the model is tasked to predict the node’s children in the graph. This task is constructed on both AST and DFG data.
We sampled about 75K graph-question-answer tuples for each (language, graph type, question type) combination, which resulted in a total of 770K GraphQA data. Further details of these GraphQA tasks and their prompts are given in Appendix B.
@@ 4.1.2 Downstream Tasks
In the main experiments, we consider the setting of multi-task finetuning (MFT), where the model is simultaneously trained on all five downstream tasks. However, we demonstrate with a small-scale experiment that GALLa can be also used for single-task finetuning (SFT).
Model Setting Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect Avg CodeGen 350M Baseline 40.2/42.3 94.4 9.0 13.6/8.8/12.5 56.9 34.7 G2C 42.7/43.3 94.6 8.5 13.9/9.5/12.8∗ 56.5 35.2 (+1%) G2C + GraphQA 50.0/45.3 93.9 8.6 14.0/9.5/12.7∗ 58.7 36.6 (+5%) StarCoder 1B Baseline 0.6/0.0 40.3 0.4 7.8/3.2/5.1 54.2 14.0 G2C 6.1/1.2 50.4 0.6 9.4/4.3/6.8 54.8 16.7 (+20%) G2C + GraphQA 6.1/1.2 65.8 0.3 9.5/4.5/7.4 56.5 18.9 (+36%) Phi-1 1.3B Baseline 73.8/53.3 93.5 10.1 14.6/11.3/13.8 61.1 41.4 G2C 67.7/64.3 94.3 10.9 15.1/11.1/13.4∗ 60.4 42.2 (+2%) G2C + GraphQA 72.0/66.3 94.9 10.2 15.2/11.7/13.6∗ 60.3 43.0 (+4%) Qwen2.5-Coder 1.5B Baseline 35.4/73.8 95.2 13.2 14.4/10.8/13.6 58.4 39.3 G2C 64.0/77.4 95.5 15.7 14.7∗/11.5/13.4 60.3 44.1 (+12%) G2C + GraphQA 65.2/76.0 95.7 15.0 15.0/11.3/14.3 61.3 44.2 (+12%) LLaMA2 7B Baseline 59.7/56.1 40.0 0.8 2.0/1.5/1.8 53.4 26.9 G2C 69.5/59.1 41.5 1.0 2.4/1.7/1.9∗ 53.5 28.8 (+7%) G2C + GraphQA 64.0/53.3 39.0 1.6 2.2/1.6∗/2.0 55.2 27.4 (+2%) LLaMA3 8B Baseline 80.5/74.8 94.7 12.1 14.2/11.2/12.3 56.7 44.6 G2C 80.5/77.0 94.9 13.4 14.0∗/11.7/11.8 56.7 45.0 (+1%) G2C + GraphQA 80.5∗/78.2 94.9 13.6 14.5/11.7/12.9 57.1 45.4 (+2%) Qwen2.5-Coder 14B Baseline 38.4/82.1 94.3 12.7 18.0/20.2/14.5 58.5 42.3 G2C 50.0/82.8 95.1 13.1 18.1∗/19.9/15.7 59.1 44.2 (+4%) G2C + GraphQA 64.0/83.4 94.4 13.8 17.9∗/20.5∗/15.9 58.0 46.0 (+9%)
@@ 4.2.1 Model
We use a DUPLEX (Ke et al., 2024) with 1024 hidden states and 7M parameters as the GNN, and a cross-attention layer with randomly initialized learnable queries as the adapter. For the LLM, we consider six distinct models of varying sizes: CodeGen 350M (Nijkamp et al., 2023), StarCoder 1B (Li et al., 2023), Phi-1 1.3B (Gunasekar et al., 2023), LLaMA2 7B (Touvron et al., 2023) LLaMA3 8B (Dubey et al., 2024), and Qwen2.5-Coder 1.5B & 14B (Hui et al., 2024). All of these models are prerained (partially) on source code data, and demonstrate strong performance on code-related downstream tasks.
@@ 4.2.2 Training
The second stage of training largely follows the same setting but with a smaller learning rate (5e-5) and a smaller global batch size (96). The model is trained for 5 epochs on the mixture of downstream task data and graph alignment data in stage 2, and the checkpoint with the lowest validation loss on the downstream tasks is chosen for evaluation. Among the different LLMs, CodeGen, StarCoder, Phi-1, and Qwen2.5-Coder 1.5B are trained in full scale, while LLaMA2, LLaMA3, and Qwen2.5-Coder 14B are trained using LoRA (Hu et al., 2022) with rank 64. All training in this stage takes place on a single machine with 8 A100s.
@@ 4.3 Results
The results of applying Graph2Code and GraphQA alignment to the code finetuning process of five models are presented in Table 3. Graph alignment brings consistent improvement over the baseline model, especially on the weaker backbones such as StarCoder, increasing the average performance on five tasks by up to 36%.
Notably, while our graph alignment data include only Python and Java programs, from Table 3 we observe that they can even improve six of the seven models’ performance on tasks in other languages - code summarization in JavaScript, and defect detection in C. This showcases that the knowledge about code structures acquired in GALLa can be generalized across programming languages, as learning to align to Python and Java structural graphs improves the finetuning performance of downstream tasks in other languages.
In Table 4, we also present the results of single-task finetuning with Phi-1. In this setting, we find that Graph2Code still improves the average performance on all tasks by 2%, and by 3% on Python and Java tasks. On the other hand, GraphQA brings limited improvement, and even a large drop on the C defect detection task. We hypothesize that this is because the diverse data in GraphQA serve a similar role to instruction tuning for LLMs (Sanh et al., 2022; Chung et al., 2022), where the benefits of cross-task transfer only start to manifest when the number of tasks is large. We find that models trained with single-task finetuning are prone to hallucinations, where they do not follow instructions in test samples but output answer templates learned from the GraphQA data instead (see Appendix D for examples).
@@
Training Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect Avg MLP G2C 73.2/66.9 95.2 11.7 15.1/11.6/14.1 61.6 43.7 G2C + GQA 76.2/68.9 95.3 12.7 14.9/12.1/13.6 59.9 44.2 MagNet G2C 75.6/65.3 94.9 11.2 14.5/11.3/13.9 63.1 43.7 G2C + GQA 74.4/65.1 95.0 12.8 15.0/12.4/14.1 63.2 44.0